# Exact treatment of the quantum Langevin equation under time-dependent system-bath coupling via a train of delta distributions

### v1

@@ II QLE with time-dependent interaction
In this section, we introduce the QLE with a time-dependent coupling. We begin by considering the Caldeira-Leggett model, where the environment is modeled as a collection of $\mathcal{N}$ independent quantum harmonic oscillators. Note that, at this stage, the environment contains a finite number of oscillators. After solving the QLE, we calculate the correlation functions and the covariance matrix of the system and then take the limit $\mathcal{N}\to\infty$.
@@
where $c(t)$ represents the time-dependent coupling between the system and the environment. For convenience, we write $c(t)=c\chi(t)$, where $c>0$ is the coupling constant and $\chi(t)$ is the switching function, which describes the time dependence of the interaction. In what follows, we assume that the interaction begins at time $t=0$.
@@
where $q_{j}^{(\text{h})}(t)$ is the homogeneous solution to the equation, and $\Theta(t)$ is Heaviside’s step function. Inserting this solution into the right-hand side of (4a) gives the QLE:
@@
where $\Sigma(t)$ is the so-called dissipation kernel (also known as the memory kernel) defined by
@@
and the environment’s operator $\zeta(t)$, known as the noise term, is given by
@@
We now comment on our derived QLE. Observe that, in general, the dissipation kernel $\Sigma(t,t^{\prime})$ is not time-translation invariant (i.e., nonstationary). An exception occurs when the switching function $\chi(t)$ is taken as a constant, $\chi(t)=1$, leading to $\Sigma(t,t^{\prime})\equiv\Sigma(t-t^{\prime})=-\Gamma(t-t^{\prime})$, so that the dissipation kernel only depends on the time difference (i.e., stationary). This particular case is widely employed in the literature, and the stationary QLE can be straightforwardly solved using the Laplace transformation. It is important to note that if the dissipation kernel is stationary, $\Sigma(t,t^{\prime})=\Sigma(t-t^{\prime})$, then the Laplace transform of the convolution integral is simply $\tilde{\Sigma}(z)\tilde{Q}(z)$ due to the convolution property. Here, $z\in\mathbb{C}$ is the Laplace variable, and $\tilde{Q}(z)\equiv\mathcal{L}[Q(t)]$, where $\mathcal{L}$ represents the Laplace transform with respect to $t$. In this case, the resulting algebraic equation in the Laplace domain leads to
@@
where $\tilde{G}(z)\coloneqq(z^{2}+\Omega^{2}+\tilde{\Sigma}(z))^{-1}$. Notice that $\zeta(t)=\xi(t)$. The inverse Laplace transformation gives us the general solution
@@
Nevertheless, if we do not assume a constant switching function, the dissipation kernel loses time-translation invariance, causing difficulties as we can no longer apply the convolution property. This is the main obstacle when considering an arbitrary switching function $\chi(t)$. In the following, we circumvent this issue by employing a collection of Dirac delta distributions.
III QLE with a train of Dirac delta
III.1 Solving the QLE
Suppose the switching function is compactly supported on $t\in[0,T]$. We choose the switching function in such a way that it is represented by $N$-Dirac delta distributions:
$\equiv\dfrac{T}{N}\sum_{k=1}^{N}\chi(t_{k})\delta(t-t_{k})\,,$ (10)
where $t_{k}\equiv kT/N$, and the Dirac deltas are uniformly distributed between $t\in[0,T]$. Note that $T/N$ is the interval between successive interactions, $t_{k}$ and $t_{k+1}$. This allows us to simplify the term with the dissipation kernel in the QLE (6). Specifically, the Laplace transform of this term reads
$\mathcal{L}\int_{0}^{t}\text{d}t^{\prime}\,\Sigma(t,t^{\prime})Q(t^{\prime})$
$=-\dfrac{T^{2}}{N^{2}}\sum_{k,l=1}^{N}\chi(t_{k})\chi(t_{l})\Gamma(t_{k}-t_{l})\Theta(t_{k}-t_{l})Q(t_{l})e^{-zt_{k}}\,.$ (11)
The noise term $\chi(t)\xi(t)$ can also be transformed as
Therefore, after applying the Laplace transformation to the QLE and obtaining $\tilde{Q}(z)$, we perform the inverse Laplace transformation to obtain
$Q(t)$ $=Q_{t}^{(0)}+\sum_{l=1}^{N}K_{t,t_{l}}Q_{l}+\Xi_{t}\,,$ (13)
where
$Q_{t}^{(0)}\coloneqq Q(0)\cos(\Omega t)+\dfrac{P(0)}{\Omega}\sin(\Omega t)\,,$ (14a)
$K_{t,t_{l}}\equiv\dfrac{T^{2}}{N^{2}}\sum_{k=1}^{N}\Sigma(t_{k},t_{l})\Theta(t_{k}-t_{l})$
×$\qquad\times\dfrac{\sin[\Omega(t-t_{k})]}{\Omega}\Theta(t-t_{k})\,,$ (14b)
$Q_{l}\equiv Q(t_{l})\,,$ (14c)
$\Xi_{t}\equiv\sum_{k=1}^{N}\zeta(t_{k})\dfrac{\sin[\Omega(t-t_{k})]}{\Omega}\Theta(t-t_{k})\,.$ (14d)
Here, $Q_{t}^{(0)}$ is the solution to the Heisenberg equation of motion for a free quantum harmonic oscillator, $K_{t,t_{l}}$ is the term that emerges from the dissipation kernel (thus, responsible for the memory effect) with the property $K_{t,t_{N}}=0$ due to $\Sigma(t_{k},t_{k})=0$, $\Xi_{t}$ corresponds to the noise term, and $Q_{l}$ is $Q(t)$ at time $t=t_{l}$. Although $Q(t)$ in (13) takes a very simple form, it still depends on itself in the past, $Q_{l}$. Below, we obtain the general solution $Q(t)$ that depends only on the initial values $Q(0)$ and $P(0)$.
To this end, we express $Q_{l}$ in terms of $Q(0)$ and $P(0)$. Substituting $t=t_{l}$ in (13) gives us
$Q_{l}$ $=Q_{l}^{(0)}+\sum_{i=1}^{N}K_{li}Q_{i}+\Xi_{l}\,,$ (15)
where $Q_{l}^{(0)}$ and $\Xi_{l}$ are understood as $Q_{l}^{(0)}\equiv Q^{(0)}(t_{l})$ and $\Xi_{l}\equiv\Xi(t_{l})$. The equation above can be written in terms of vectors and matrices as follows:
$\bm{Q}$ $=\bm{Q}^{(0)}+\mathbf{K}\bm{Q}+\bm{\Xi}\,,$ (16)
where
$\bm{Q}$ $\coloneqq[Q_{1},Q_{2},\ldots,Q_{N}]^{\intercal}\,,$ (17a)
$\bm{Q}^{(0)}$ $\coloneqq[Q^{(0)}_{1},Q^{(0)}_{2},\ldots,Q^{(0)}_{N}]^{\intercal}\,,$ (17b)
$\mathbf{K}$ $\coloneqq\begin{bmatrix}0&0&0&\ldots&\ldots&0\\ 0&0&0&\ldots&\ldots&0\\ K_{31}&0&0&\ldots&\ldots&0\\ K_{41}&K_{42}&\ddots&\ddots&&\vdots\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ K_{N1}&K_{N2}&\ldots&K_{N,N-2}&0&0\\ \end{bmatrix}\,,$ (17c)
$\bm{\Xi}$ $\coloneqq[\Xi_{1},\Xi_{2},\ldots,\Xi_{N}]^{\intercal}\,.$ (17d)
Here, $K_{li}$ is the elements of the $N\times N$ matrix $\mathbf{K}$ given by
$K_{li}$ $=\dfrac{T^{2}}{N^{2}}\sum_{k=1}^{N}\Sigma(t_{k},t_{i})\Theta(t_{k}-t_{i})\dfrac{\sin[\Omega(t_{l}-t_{k})]}{\Omega}\Theta(t_{l}-t_{k})\,,$ (18)
and the matrix $\mathbf{K}$ is a strictly lower triangular matrix due to Heaviside’s step functions in $K_{li}$. Note that $K_{i+1,i}=0$ since $\Sigma(t_{k},t_{k})=0$.
The equation (16) can be solved for $\bm{Q}$ as
$\bm{Q}$ $=(\mathbf{I}-\mathbf{K})^{-1}\bm{Q}^{(0)}+(\mathbf{I}-\mathbf{K})^{-1}\bm{\Xi}\,,$ (19)
because
$=\mathbf{I}-\mathbf{K}^{N}=\mathbf{I}-\mathbf{O}=\mathbf{I}\,.$ (21)
We thus obtained a compact form of the vector of $Q_{l}$:
$\bm{Q}$ $=\mathsf{K}\bm{Q}^{(0)}+\mathsf{K}\bm{\Xi}\,.$ (22)
Substituting $Q_{l}$ into $Q(t)$ in (13), we finally reached our main result:
@@ III.1 Solving the QLE
The function $G[f](t)$ in (24) explicitly shows how the memory effects of the environment affect the system. If the dissipation kernel is negligible, then $K_{t,t_{l}}$ and $\mathsf{K}_{li}$ vanish. In what follows, we denote $G_{f}(t)\equiv G[f](t)$ for brevity.
Equation (23) is the exact solution to the QLE (6) when a compactly supported switching function is described by the train of Dirac deltas (10). Note that this solution is also valid for intermediate times $t\in(0,T)$, as the Heaviside step functions naturally eliminate irrelevant terms. Moreover, at each discrete time $t_{l}$, $l\in\{1,2,\ldots,N\}$, the solution $Q(t=t_{l})$ reduces to Eq. (22), which can be expressed using a single function $\mathcal{G}(t)$ as
@@
III.2 Pictorial interpretation
Each term in our result (23) consists of time-dependent functions $G_{f_{Q}}(t),G_{f_{P}}(t)$, and $G_{f_{\zeta}}(t)$, which are given by (24). These functions are written in the form
$G_{f}(t)$ $=f_{t}+\sum_{l,i=1}^{N}K_{t,t_{l}}\mathsf{K}_{li}f_{i}\,,$ (27)
where $f_{t}\equiv f(t)$. This expression clearly shows the influence of the dissipation kernel, as both $K_{t,t_{l}}$ and $\mathsf{K}_{li}$ contain $\Sigma(t)$. Although the second term appears cumbersome, we show that it can be intuitively understood using diagrams.
Figure 2: The diagrams illustrating $G_{f}(t)$ in (27) for (a) $N=2$, (b) $N=3$, and (c) $N=4$. A loop connecting $t_{i}$ and $t_{j}$ ($j>i$) corresponds to $\Sigma(t_{j},t_{i})\equiv\Sigma_{ji}$.
To begin with, we introduce diagrams that characterize each term in (27). The function $f_{t}$ describes $G_{f}(t)$ in the absence of the dissipation kernel. We depict $f_{t}$ as a dashed line along the time axis, as shown in Fig. 1(i). In particular, $f_{i}\equiv f(t_{i})$ is illustrated as a dashed line from $t=0$ to $t=t_{i}$ [Fig. 1(ii)]. The quantity $K_{t,t_{l}}$, defined in (14b), is the sum of the products of $\Sigma(t_{k},t_{l})\Theta(t_{k}-t_{l})$ and $\sin[\Omega(t-t_{k})]\Theta(t-t_{k})/\Omega$. The factor $\Sigma(t_{k},t_{l})\Theta(t_{k}-t_{l})$ is interpreted as an “exchange” of a signal from $t=t_{l}$ to $t_{k}$ due to the environment’s memory effect. Subsequently, the term $\sin[\Omega(t-t_{k})]\Theta(t-t_{k})/\Omega$ propagates this effect from $t=t_{k}$ to $t(>T)$. Thus, we assign a solid loop to $\Sigma(t_{k},t_{l})$ and a line to $\sin[\Omega(t-t_{k})]\Theta(t-t_{k})/\Omega$ as illustrated in Fig. 1(iii). A similar interpretation applies to $K_{li}$ defined in (18), except that the propagation terminates at $t=t_{l}$ [Fig. 1(iv)].
We note that $\Sigma(t_{k},t_{k})=0$ for any environment. This means that an “instantaneous loop”, which is closed at a single time $t=t_{k}$, is not allowed. Physically, it guarantees that there is no immediate backreaction from the environment to the system.
Let us use these diagrams to understand the role of memory effects. We first note that the elements of the matrix $\mathsf{K}$ consist of $K_{li}$. For instance,
$\mathsf{K}$ $=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\quad\text{for }N=2\,,$
$\mathsf{K}$ $=\begin{bmatrix}1&0&0\\ 0&1&0\\ K_{31}&0&1\end{bmatrix}\quad\text{for }N=3\,,$
$\mathsf{K}$ $=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ K_{31}&0&1&0\\ K_{41}&K_{42}&0&1\end{bmatrix}\quad\text{for }N=4\,,$
$\mathsf{K}$ $=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ K_{31}&0&1&0&0\\ K_{41}&K_{42}&0&1&0\\ K_{51}+K_{53}K_{31}&K_{52}&K_{53}&0&1\end{bmatrix}\quad\text{for }N=5\,.$
Let us take a simple example of $N=2$ and visualize $G_{f}(t)$ in (27). Reminding that $K_{t,t_{N}}=0$, the function $G_{f}(t)$ for $N=2$ can be explicitly written as
$G_{f}(t)$ $=f_{t}+\sum_{l,i=1}^{N=2}K_{t,t_{l}}\mathsf{K}_{li}f_{i}$
$=f_{t}+\begin{bmatrix}K_{t,t_{1}}&0\end{bmatrix}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{bmatrix}f_{1}\\ f_{2}\end{bmatrix}$
$=f_{t}+K_{t,t_{1}}f_{1}\,.$ (28)
Figure 2(a) illustrates this expression using our diagrammatic elements. The term $K_{t,t_{1}}f_{1}$ can be understood as follows. First, draw a dashed line representing $f_{1}$ from $t=0$ to $t_{1}$. Then, we consider the factor $K_{t,t_{1}}$, which contains a loop and a line as illustrated in Fig. 1(iii). In general, all possible configurations must be considered for each term. In the $N=2$ case, there is only one configuration for $K_{t,t_{1}}f_{1}$ as shown in Fig. 2(a). For $N=3$, however, there are two possible configurations for $K_{t,t_{1}}f_{1}$ as depicted in Fig. 2(b), and their sum gives the final value of $K_{t,t_{1}}f_{1}$.
For $N=4$, various configurations emerge [Fig. 2(c)]. The same argument above applies to $K_{t,t_{2}}f_{2}$ and $K_{t,t_{3}}f_{3}$. We also have the term $K_{t,t_{3}}K_{31}f_{1}$, which contains two loops arising from $K_{t,t_{3}}$ and $K_{31}$. Here, $K_{31}$ is represented by a loop starting from $t_{1}$ and a line ending at $t_{3}$.
Overall, the expression for $G_{f}(t)$ sums over all possible memory effects of the environment. We stress that both Markovian and non-Markovian effects can also be understood using the diagrams. Markovian processes correspond to diagrams that consist of a single loop connecting adjacent times, namely, the terms with $\Sigma(t_{i+1},t_{i})$. One can be convinced by considering the “continuum limit” of the delta switchings, $N\to\infty$ and $t_{i+1}-t_{i}(\equiv T/N)\to 0$. In this case, the adjacent memory effect $\Sigma(t_{i+1},t_{i})$ can be considered time-local. On the other hand, non-Markovian effects are represented by loops connecting distant points $t_{i}$ and $t_{i+j}$ with $j\geq 2$. In Fig. 2(c), the Markovian diagrams correspond to $\Sigma_{21}$, $\Sigma_{32}$, and $\Sigma_{43}$ in $K_{t,t_{1}}f_{1}$, $K_{t,t_{2}}f_{2}$, and $K_{t,t_{3}}f_{3}$, respectively. Note that the diagram corresponding to $K_{t,t_{3}}K_{31}f_{1}$, which contains two adjacent loops, is not interpreted as Markovian. The mathematical argument is provided in Appendix B.
III.3 Two-time correlation functions and the covariance matrix
@@
III.3.1 Two-time correlation functions
@@
III.3.2 The covariance matrix
In continuous-variable quantum mechanics, a quantum state is called Gaussian if it is determined solely by the first and second moments. We denote the first moment by $\bar{\bm{R}}\equiv\braket{\bm{R}}$. For the second moments, one typically considers the covariance matrix—a real symmetric, positive-definite matrix—$\mathbb{V}(t)$ for describing the second moments, and it is given by [21]
@@ III.3.2 The covariance matrix
A particularly useful tool for studying the covariance matrix in Gaussian systems is the Gaussian operations. These are the completely-positive (CP) maps that transform Gaussian states to Gaussian states [20]. To illustrate this, suppose a joint system is initially prepared in a product of Gaussian states, $\rho_{\text{tot}}(0)=\rho_{\text{s}}(0)\otimes\rho_{\text{E}}(0)$, which is also Gaussian. A unitary time-evolution operator generated by a Hamiltonian composed of a linear and a quadratic term in $R_{i}$ is an example of the Gaussian operation known as the Gaussian unitary transformation [20]. The Hamiltonian in the Caldeira-Leggett model used in this paper indeed generates such a unitary transform, and this fact does not change when a train of delta-switchings is employed. Moreover, a partial trace of a Gaussian joint state is also a Gaussian operation. Thus, the system’s reduced quantum state $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U(t)\rho_{\text{tot}}(0)U^{\dagger}(t)]$ after the time-evolution remains to be Gaussian. This fact can be utilized to characterize the evolution of the statistical moments. Assuming the initial joint state is separable Gaussian and they evolve under a trace-preserving Gaussian operation (hence, it is a CPTP map), the first moment and the covariance matrix evolve as [20]
@@
where $\bm{\Omega}$ is the symplectic form that appears in the canonical commutation relations: $[R_{i},R_{j}]=\mathsf{i}\Omega_{ij}$.
@@
where $\mathbb{T}_{t}$ and $\mathbb{N}_{t}$ are the matrices associated with $\mathcal{E}_{t,0}$. We prove these properties in Appendix A.
@@
$N_{QQ}(t)$ $=\sum_{k,k^{\prime}=1}^{N}G_{f_{\zeta}^{(k)}}(t)G_{f_{\zeta}^{(k^{\prime})}}(t)\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,,$ (38a)
$N_{QP}(t)$ $=\sum_{k,k^{\prime}=1}^{N}G_{f_{\zeta}^{(k)}}(t)\dot{G}_{f_{\zeta}^{(k^{\prime})}}(t)\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,,$ (38b)
$N_{PQ}(t)$ $=\sum_{k,k^{\prime}=1}^{N}\dot{G}_{f_{\zeta}^{(k)}}(t)G_{f_{\zeta}^{(k^{\prime})}}(t)\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,,$ (38c)
$N_{PP}(t)$ $=\sum_{k,k^{\prime}=1}^{N}\dot{G}_{f_{\zeta}^{(k)}}(t)\dot{G}_{f_{\zeta}^{(k^{\prime})}}(t)\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,.$ (38d)
It turns out that the matrices $\mathbb{T}_{t}$ and $\mathbb{N}_{t}$ for the case of delta-switchings do not satisfy the Markovianity conditions (35) in general. However, one can reduce the dynamics to Markovian by applying the Born-Markov approximation. We demonstrate this fact in Appendix B for an environment with the so-called Lorentz-Drude spectral density described in the next section.
IV Demonstration with the Lorentz-Drude spectral density
IV.1 Spectral density
Our main result in (23) was derived under the assumption that the environment consists of a finite collection of quantum harmonic oscillators. Nevertheless, we can extend this result to the continuum limit by introducing the spectral density $\sigma(\omega)$ of the environment. In the traditional QLE with a continuous constant coupling, the spectral density is defined via the Fourier transform of the dissipation kernel. In contrast, we illustrate that the spectral density should be defined by using the discrete-time Fourier transform (DTFT) when a train of delta switchings is employed.
IV.1.1 Introducing the spectral density
Let us introduce the traditional spectral density used when a continuous constant coupling is considered. Our first aim is to rewrite the quantities that appear in $\braket{R_{i}(t)}$ and $\braket{R_{i}(t)R_{j}(t^{\prime})}$ in terms of the spectral density $\sigma(\omega)$. Consider the time-translation invariant part $\Gamma(t)$ of the dissipation kernel. For a finite collection of quantum harmonic oscillators, $\Gamma(t)$ is given by (7b). We then introduce the spectral density $\sigma(\omega)$ by Fourier transforming $\Gamma(t)$ as [^3] In this paper, the Fourier transform is denoted by $\hat{\Gamma}(\omega)\equiv\mathcal{F}[\Gamma(t)]$. This notation should not be confused with that used for linear operators in quantum mechanics.

### v3

@@ IV.1 Solving the QLE
We begin by considering the Caldeira–Leggett model, where the environment is modeled as a collection of $\mathcal{N}$ independent quantum harmonic oscillators. Note that, at this stage, the environment contains a finite number of oscillators. After solving the QLE, we calculate the correlation functions and the covariance matrix of the system and then take the limit $\mathcal{N}\to\infty$.
@@
where $c(t)$ represents the time-dependent coupling between the system and the environment.
@@
where $q_{j}^{(\text{h})}(t)$ is the homogeneous solution to the equation. Inserting this solution into the right-hand side of (49a) gives the QLE:
@@
where the memory kernel $\Sigma(t)$ is defined by [^4] For the QLE, we will adopt the convention that a minus appears in front of $\Gamma(t)$.
@@
and the noise term is given by
@@
As explained in Sec. II.1, the integro-differential equation can be solved when the memory kernel is stationary. In this case, $\Gamma(t,t^{\prime})$ is already stationary, so $\Sigma(t,t^{\prime})$ becomes time-translation invariant if the system constantly couples to the environment: $\chi(t)=1$.
In this scenario, the resulting algebraic equation in the Laplace domain leads to
@@
where $\tilde{G}(z)\coloneqq(z^{2}+\Omega^{2}+\tilde{\Sigma}(z))^{-1}$. Notice that $\zeta(t)=\xi(t)$. The inverse Laplace transformation gives us the well-known general solution
@@
We now consider a general switching function $\chi(t)$ and apply our result (29). The free solution $Q^{(0)}_{t}$ and the free Green function $G_{0}(t)$ for a quantum harmonic oscillator are given by
$Q^{(0)}_{t}=Q(0)\cos(\Omega t)+\dfrac{P(0)}{\Omega}\sin(\Omega t)\,$,
$G_{0}(t)=\dfrac{\sin(\Omega t)}{\Omega}\,$,
and thus Eq. (29) with these functions is the exact solution to the QLE (51) when a compactly supported switching function is described by the train of Dirac deltas (15). For the sake of later use, we write the solution to the QLE in the following form:
@@
In what follows, we denote $G_{f}(t)\equiv G[f](t)$ for brevity.
Equation (55) is the exact solution to the QLE (51) when a compactly supported switching function is described by the train of Dirac deltas (15). Note that this solution is also valid for intermediate times $t\in(0,T)$, as the Heaviside step functions naturally eliminate irrelevant terms. Moreover, at each discrete time $t_{l}$ for $l\in\{1,2,\ldots,N\}$, the solution $Q(t=t_{l})$ reduces to Eq. (28), which can be expressed using a single function $\mathcal{G}(t)$ as
@@
IV.2 Two-time correlation functions and the covariance matrix
@@
IV.2.1 Two-time correlation functions
@@
IV.2.2 The covariance matrix
In continuous-variable quantum mechanics, a quantum state is called Gaussian if it is determined solely by the first and second moments. We denote the first moment by $\bar{\bm{R}}\equiv\braket{\bm{R}}$. For the second moments, one typically considers the covariance matrix $\mathbb{V}(t)$—a real symmetric, positive-definite matrix—for describing the second moments, and it is given by [1]
@@ IV.2.2 The covariance matrix
A particularly useful tool for studying the covariance matrix in Gaussian systems is the Gaussian operations. These are the completely-positive (CP) maps that transform Gaussian states to Gaussian states [28]. To illustrate this, suppose a joint system is initially prepared in a product of Gaussian states, $\rho_{\text{tot}}(0)=\rho_{\text{s}}(0)\otimes\rho_{\text{E}}(0)$, which is also Gaussian. A unitary time-evolution operator generated by a Hamiltonian composed of a linear and a quadratic term in $R_{i}$ is an example of the Gaussian operation known as the Gaussian unitary transformation [28]. The Hamiltonian in the Caldeira–Leggett model used in this paper indeed generates such a unitary transform, and this fact does not change when a train of delta-switchings is employed. Moreover, a partial trace of a Gaussian joint state is also a Gaussian operation. Thus, the system’s reduced quantum state $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U(t)\rho_{\text{tot}}(0)U^{\dagger}(t)]$ after the time-evolution remains to be Gaussian. This fact can be utilized to characterize the evolution of the statistical moments. Assuming the initial joint state is separable Gaussian and they evolve under a trace-preserving Gaussian operation (hence, it is a CPTP map), the first moment and the covariance matrix evolve as [28]
@@
where $\bm{\Omega}$ is the symplectic form that appears in the canonical commutation relations: $R_{i},R_{j}]=\mathsf{i}\Omega_{ij}$.
@@
where $\mathbb{T}_{t}$ and $\mathbb{N}_{t}$ are the matrices associated with $\mathcal{E}_{t,0}$. We prove these properties in Appendix D.
@@
$N_{QQ}(t)$ $=\sum_{k,k^{\prime}=1}^{N}G_{f_{\zeta}^{(k)}}(t)G_{f_{\zeta}^{(k^{\prime})}}(t)\nu_{k,k^{\prime}}\,$, (68a)
$N_{QP}(t)$ $=\sum_{k,k^{\prime}=1}^{N}G_{f_{\zeta}^{(k)}}(t)\dot{G}_{f_{\zeta}^{(k^{\prime})}}(t)\nu_{k,k^{\prime}}\,$, (68b)
$N_{PQ}(t)$ $=\sum_{k,k^{\prime}=1}^{N}\dot{G}_{f_{\zeta}^{(k)}}(t)G_{f_{\zeta}^{(k^{\prime})}}(t)\nu_{k,k^{\prime}}\,$, (68c)
$N_{PP}(t)$ $=\sum_{k,k^{\prime}=1}^{N}\dot{G}_{f_{\zeta}^{(k)}}(t)\dot{G}_{f_{\zeta}^{(k^{\prime})}}(t)\nu_{k,k^{\prime}}\,$, (68d)
and $\nu_{k,k^{\prime}}\equiv\braket{\{\zeta(t_{k}),\zeta(t_{k^{\prime}})\}}$. It turns out that the matrices $\mathbb{T}_{t}$ and $\mathbb{N}_{t}$ for the case of delta-switchings do not satisfy the Markovianity conditions (65) in general. However, one can reduce the dynamics to Markovian by applying the Born–Markov approximation. We demonstrate this fact in Appendix E for an environment with the so-called Lorentz-Drude spectral density described in the next section.
IV.3 Demonstration with the Lorentz–Drude spectral density
Our solution to the nonstationary QLE in (55) was derived under the assumption that the environment consists of a finite collection of quantum harmonic oscillators. Nevertheless, we can extend this result to the continuum limit by introducing the spectral density $\sigma(\omega)$ of the environment. In the traditional QLE with a continuous constant coupling, the spectral density is defined via the Fourier transform of the memory kernel. In contrast, we illustrate that the spectral density should be defined by using the discrete-time Fourier transform (DTFT) when a train of delta switchings is employed.
IV.3.1 Introducing the spectral density
Let us introduce the traditional spectral density used when a continuous constant coupling is considered. Our first aim is to rewrite the quantities that appear in $\braket{R_{i}(t)}$ and $\braket{R_{i}(t)R_{j}(t^{\prime})}$ in terms of the spectral density $\sigma(\omega)$. Consider the time-translation invariant part $\Gamma(t)$ of the memory kernel. For a finite collection of quantum harmonic oscillators, $\Gamma(t)$ is given by (52b). We then introduce the spectral density $\sigma(\omega)$ by Fourier transforming $\Gamma(t)$ as [^5] In this paper, the Fourier transform is denoted by $\hat{\Gamma}(\omega)\equiv\mathcal{F}[\Gamma(t)]$. This notation should not be confused with that used for linear operators in quantum mechanics.